# Amortized Variational Transdimensional Inference

### v1

@@ 3 Formulating a model weights distribution
Formulating and estimating $q_{\psi}$ is not as straightforward as that of $q_{\phi}$ because the discrete random variables $m\sim q_{\psi}$ are not automatically linked to the density parameters $\psi$ by automatic differentiation. This problem naturally lends itself to methods developed in black-box variational inference (Ranganath et al., , 2014; Titsias and Lázaro-Gredilla, , 2014; Wingate and Weber, , 2013) and multi-armed bandits (Bubeck and Cesa-Bianchi, , 2012), as described below. The representation of $m$ is any discrete random variable on a finite space $\pazocal{M}$. Writing the true distribution of $m$ as $\pi_{m}$, a finite $\pazocal{M}$ implies the existence of a categorical distribution $\pi_{\zeta}$ which is bijectively equivalent to $\pi_{m}$. The random variables $\zeta\sim\pi_{\zeta}$ exist on the finite support $\zeta\in\pazocal{C}\subset\mathbb{N}$, thus $|\pazocal{C}|=|\pazocal{M}|$. This property is used by the surrogate-based approach described in Section 3.1. We formalize this concept via the following.
Proposition 3.1.
Every finite discrete distribution over a finite support $\pazocal{M}=\{m_{1},m_{2},\ldots,m_{k}\}$ has a unique representation as a categorical distribution. Specifically, there exists a bijective mapping between the set of all finite discrete distributions on $\pazocal{M}$ and the set of categorical distributions parameterized by probability vectors $\bm{\psi^{\zeta}}$ over $\pazocal{M}$. (Proof in Appendix C).
@@ 3.1 Estimation via surrogate
The objective in Equation˜10 can be rewritten as a single-variable objective with respect to $\phi$:
@@
Computing the expression above within an optimization loop over $\phi$ in practice would, however, require the evaluation of flow-based densities over the entire model space. We may, instead, follow a cheaper-to-evaluate density $q_{u,\phi}$ which approximates $q_{\ell,\phi}^{*}$ for a given $\phi$, by means of learning a surrogate model over $\ell$ within the same optimization loop [^1] We are here assuming that the prior $p(m)$ is cheap to evaluate. If not, we can model $\ell(m;\phi)+\log p(m)$, instead, with a surrogate, which leads to similar theoretical guarantees after minimal adjustments.. In particular, we derive a Gaussian process (GP) upper confidence bound (Srinivas et al., , 2010), which provides the following approximation to the optimal model probabilities:
@@
where $u_{t}(m):=\mu_{t}(m,\phi_{t})+\beta\sigma_{t}(m,\phi_{t})$, with $\mu_{t}$ and $\sigma_{t}^{2}$ representing the posterior mean and variance of a GP model conditioned on all mini-batches of data $\pazocal{B}_{t}:=\{\phi_{t-1},m_{t,i},\log h_{\phi_{t-1}}(\bm{z}_{t,i}|m_{t,i})\}_{i=1}^{B}$ available at iteration $t$ of stochastic gradient descent, and $\phi_{t}$ denotes the current flow parameters. In this form, $u_{t}$ provides an upper confidence bound (UCB) over $\ell(m;\phi_{t})$ determined by the choice of confidence parameter $\beta>0$. The GP posterior mean and variance can be derived in closed form if the observation noise is Gaussian with, e.g., variance $\sigma_{\epsilon}^{2}$. We, however, show that a sub-Gaussian noise assumption is sufficient to use a conventional GP model. In addition, if $\phi_{t}$ follows a convergent sequence (e.g., by ensuring diminishing step sizes during gradient-based optimization), we provide the following guarantee.
@@
$D_{\mathrm{KL}}(q_{u,t}||q_{\ell,\phi}^{*})\in\pazocal{O}_{\operatorname*{\mathbb{P}}}(t^{-1/2}),$ (15)
@@
This is a direct application of Theorem B.3, derived in the Appendix. The result above tells us that the UCB-based models distribution approaches the optimal distribution at a rate of $\pazocal{O}_{\operatorname*{\mathbb{P}}}(t^{-1/2})$ and ultimately converges to it as $t\to\infty$. Therefore, a stochastic gradient optimizer using samples from the surrogate density $q_{u,t}$ should asymptotically converge to the optimization path determined by the optimal $q_{\ell,\phi_{t}}^{*}$. That is, under appropriate settings for, e.g., its learning rate schedule, the optimization will converge to $\phi^{*}$.
Due to the reliance on GP-based approximations, a naive implementation of this approach would incur a cost of $\pazocal{O}(B^{3}t^{3})$ per stochastic gradient step, where $B$ is the mini-batch size due to the requirement of performing matrix inversions (Rasmussen and Williams, , 2006). However, sparse approximations to GPs can significantly reduce this cost to make it practically implementable (Rasmussen and Williams, , 2006; Gijsberts and Metta, , 2013). In particular, for our purposes, we implemented a diagonal Gaussian approximation, which makes the cost linear in the batch size and constant in $t$ via a mean-field approximation.
@@ 3.2 Categorical and neural probability mass functions
In general, by Proposition˜3.1, we may represent probability distributions over the model space $\pazocal{M}$ arbitrarily via parametrizations of categorical distributions. A drawback of the surrogate-based approach above is the need to maintain and update estimates over the entire model space, which can be impractical for spaces of very large cardinality, as one may face with DAGs, for example. Hence, we introduce two approaches based on parametric models.
Categorical model:
Assume $|\pazocal{M}|=M\in{\mathbb{N}}$. Then, for $\psi\in{\mathbb{R}}^{M}$, the distribution over $\pazocal{M}$ is defined by $q_{\psi}(m):=(\sum_{j=1}^{M}\exp\psi_{j})^{-1}\sum_{i=1}^{M}\mathbb{I}[m_{i}=m]\exp\psi_{i}\,$. The logit weights vector $\psi$ is unconstrained in ${\mathbb{R}}^{M}$ and can be jointly optimized with $\phi$ by gradient methods. Note that density evaluations and the entropy can be readily computed, although the memory required is $\pazocal{O}(|\pazocal{M}|)$.
Autoregressive model:
If the model space is too large to keep track of a categorical weights vector, another approach is to use a more structured sample generation process which allows for the number of parameters to be smaller than the dimensionality of the space, i.e., $\dim(\psi)<M$. For instance, Germain et al., (2015) proposed an autoregressive parametrization for distributions over binary strings $\bm{s}\in\{0,1\}^{d_{s}}$ via the decomposition $p_{\psi}(\bm{s})=\prod_{i=1}^{d_{s}}p_{\psi}(s_{i}|s_{1},\dots,s_{i-1})$. For each $\bm{s}$, we may assign a unique $m\in\pazocal{M}$ and therefore define $q_{\psi}(m):=p_{\psi}(\bm{s}(m))$. The conditional densities and sampling can be implemented via MADE, allowing us to map the entire model space with fewer parameters as long as $2^{d_{s}}\geq|\pazocal{M}|$. The same reasoning can be applied to a DAG via decomposition of its adjacency matrix. More details are provided in Section˜F.8 with the implementation for DAGs in Section˜F.7.
@@ 3.3 Estimation via Monte Carlo gradients
When $|\pazocal{M}|$ is too large to use a surrogate-based approach, or to even parameterize an entire vector of categorical weights in physical memory, we can employ neural-based methods that use gradient descent and estimation of the gradients of $\psi$ via Monte Carlo estimation of gradients (MCG) (Mohamed et al., , 2020). Expressed in terms of the distribution of models and its parameters $\psi$, we have
$\nabla_{\psi}q_{\psi}(m)=q_{\psi}(m)\nabla_{\psi}\log q_{\psi}(m).$ (16)
The gradient of the expectation in equation 10 with respect to $\psi$ is
@@
In practice, the variance of this estimator can be very high. However, techniques exist to reduce this variance (Paisley et al., , 2012; Ranganath et al., , 2014; Mohamed et al., , 2020) for general applications. We use a control variate $\varsigma$ in the form
@@
where
$g(\phi,\psi,\varsigma)$ $=\mathbb{E}_{\bm{z}\sim\nu_{d_{\mathrm{max}}}}\left[\log h_{\phi}(\bm{z}|m)+\log q_{\psi}(m)-\log p(m)-\varsigma\right].$ (19)
We compute $\varsigma$ using the method described in Section˜D.1 (full description in Appendix˜D).
As mentioned earlier, the benefit of using MCG for variational parameter estimation is the flexibility of choice for $q_{\psi}$. In this paper, we compare two choices: (1) MCG of the logits of a standard categorical distribution, and (2) MCG of multi-layer perceptron weights that parameterise a configuration of the MADE neural autoregressive density estimator of Germain et al., (2015) (see Section˜F.7). When $|\pazocal{M}|$ is large, such implementations of $q_{\psi}$ permit an efficient approximate representation of the true model distribution.
3.4 Controlling the optimization
An issue of practical importance is that the convergence of $\psi\rightarrow\psi^{*}$ is dependent on the convergence of $\phi\rightarrow\phi^{*}$. Conversely, optimal sample-efficiency for the inference of $\phi$ is achieved when $\psi\approx\psi^{*}$. This is intuitive because $q_{\phi}$ should “focus” primarily on the highest probability models such that variability in the approximation is minimized efficiently overall. Therefore, it is of some importance to control the variance of the estimates for $q_{\psi}$ as discussed above. For our implementation, bounding the information gain on the models categorical distribution $q_{\psi_{t}}\to q_{\psi_{t+1}}$ provided a practical way to stabilize the optimization process, similar to successful approaches in the reinforcement learning literature (Schulman et al., , 2017). We provide further details on our implementation in Section˜D.2.
@@ 4 Related work
Conditional normalizing flows (Winkler et al., , 2019; Durkan et al., , 2019) have emerged as powerful tools for incorporating conditioning information. Existing methods use the context variable as a conditioning input, but fewer adapt the flow architecture itself. An exception is the transport‐based reversible jump MCMC method (Davies et al., , 2023), which learns proposals for transdimensional moves, but does not readily allow its use as an inverse autoregressive flow (Kingma et al., , 2016). In contrast, we introduce an identity‐parameterized CoSMIC transformation without identity-map training. We bypass path-wise approximations to discrete distributions (Jang et al., , 2017; Maddison et al., , 2017), instead comparing Monte Carlo gradient estimation (Mohamed et al., , 2020) with Bayesian optimization (Shahriari et al., , 2016). We adopt an information-based approach to scale gradient steps using “small steps,” inspired by reinforcement learning (Schulman et al., , 2017). Bayesian methods for model selection and optimization have advanced with black‐box variational inference (Ranganath et al., , 2014; Titsias and Lázaro-Gredilla, , 2014; Wingate and Weber, , 2013) and flexible flows (Rezende and Mohamed, , 2015; Papamakarios et al., , 2017; Durkan et al., , 2019). Recent work in amortized Bayesian mixture models (Kucharsky and Burkner, , 2025) shows amortization over multiple mixture components using conditional normalizing flows, but not for variable dimensions. Conversely, Li et al., (2020) introduces an architecture for learning imputation over transdimensional inputs, but lacks immediate application as a variational density. Our approach unifies transdimensional inference with flow‐based variational methods, bypassing the need for tailored dimension jumps and broadening applications.
@@ 5 Experiments
We present experiments involving synthetic and real data on two representative applications: robust variable selection and directed acyclic graphs. To evaluate the quality of the approximation $q_{\psi,\phi}(m,\bm{\theta}_{m})$ to the target distribution $\pi(m,\bm{\theta}_{m})$ for a relatively small $|\pazocal{M}|<2^{19}$ model space, we use the average negative log-likelihood (NLL) computed over a set of samples drawn from $\pi$ via a baseline sampling method, in this case reversible jump MCMC (Sisson, , 2005). Let $\{(m^{i},\bm{\theta}_{m}^{i})\}_{i=1}^{N}$ denote $N$ independent samples from $\pi(m,\bm{\theta}_{m})$. The average NLL corresponds to the cross-entropy $H(\pi,q_{\psi,\phi})$ between the distributions $\pi$ and $q_{\psi,\phi}$, which quantifies the expected number of bits needed to encode samples from $\pi$ using the distribution $q_{\psi,\phi}$, and is defined as $\text{NLL}=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi,\phi}(m^{i},\bm{\theta}_{m}^{i})$. Standard metrics Kummerfeld and Rix, (2019) can compare VTI DAG inference quality with baseline frequentist approaches.

### v2

@@ 3 Formulating a model weights distribution
Formulating and estimating $q_{\psi}$ is not as straightforward as that of $q_{\phi}$ because the discrete random variables $m\sim q_{\psi}$ are not automatically linked to the density parameters $\psi$ by automatic differentiation. This problem naturally lends itself to methods developed in black-box variational inference [Ranganath et al., 2014, Titsias and Lázaro-Gredilla, 2014, Wingate and Weber, 2013] and multi-armed bandits [Bubeck and Cesa-Bianchi, 2012], as described below. The representation of $m$ is any discrete random variable on a finite space $\pazocal{M}$. Writing the true distribution of $m$ as $\pi_{m}$, a finite $\pazocal{M}$ implies the existence of a categorical distribution $\pi_{\zeta}$ which is bijectively equivalent to $\pi_{m}$. The random variables $\zeta\sim\pi_{\zeta}$ exist on the finite support $\zeta\in\pazocal{C}\subset\mathbb{N}$, thus $|\pazocal{C}|=|\pazocal{M}|$. This property is used by the surrogate-based approach described in Section 3.1.We formalize this concept via Proposition D.1.
@@ 3.1 Estimation via surrogate
The objective in Equation 9 can be rewritten as a single-variable objective with respect to $\phi$:
@@
Computing the expression above within an optimization loop over $\phi$ in practice would, however, require the evaluation of flow-based densities over the entire model space. We may, instead, follow a cheaper-to-evaluate density $q_{u,\phi}$ which approximates $q_{\ell,\phi}^{*}$ for a given $\phi$, by means of learning a surrogate model over $\ell$ within the same optimization loop [^2] We are here assuming that the prior $p(m)$ is cheap to evaluate. If not, we can model $\ell(m;\phi)+\log p(m)$, instead, with a surrogate, which leads to similar theoretical guarantees after minimal adjustments.. In particular, we derive a Gaussian process (GP) upper confidence bound [Srinivas et al., 2010], which provides the following approximation to the optimal model probabilities:
@@
where $u_{t}(m):=\mu_{t}(m,\phi_{t})+\beta\sigma_{t}(m,\phi_{t})$, with $\mu_{t}$ and $\sigma_{t}^{2}$ representing the posterior mean and variance of a GP model conditioned on all mini-batches of data $\pazocal{B}_{t}:=\{\phi_{t-1},m_{t,i},\log h_{\phi_{t-1}}(\bm{z}_{t,i}|m_{t,i})\}_{i=1}^{B}$ available at iteration $t$ of stochastic gradient descent, and $\phi_{t}$ denotes the current flow parameters. In this form, $u_{t}$ provides an upper confidence bound (UCB) over $\ell(m;\phi_{t})$ determined by the choice of confidence parameter $\beta\geq 0$. The GP posterior mean and variance can be derived in closed form if the observation noise is Gaussian with, e.g., variance $\sigma_{\epsilon}^{2}$. We, however, show that a sub-Gaussian noise assumption is sufficient to use a conventional GP model. In addition, if $\phi_{t}$ follows a convergent sequence (e.g., by ensuring diminishing step sizes during gradient-based optimization), we have the following guarantee.
@@
$D_{\mathrm{KL}}(q_{u,t}||q_{\ell,\phi_{t}}^{*})\in\pazocal{O}_{\operatorname*{\mathbb{P}}}(t^{-1/2}),$ (13)
@@
The result above tells us that the UCB-based models distribution approaches the optimal distribution at a rate of $\pazocal{O}_{\operatorname*{\mathbb{P}}}(t^{-1/2})$ and ultimately converges to it as $t\to\infty$. Therefore, a stochastic gradient optimizer using samples from the surrogate density $q_{u,t}$ should asymptotically converge to the optimization path determined by the optimal $q_{\ell,\phi_{t}}^{*}$. That is, under appropriate settings for, e.g., its learning rate schedule, the optimization will converge to $\phi^{*}$. Lastly, note that the result in 3.1 is independent of the choice of $\beta$, which can be set to $\beta=0$. Our analysis is mainly based on obtaining enough samples almost everywhere across the model space, which can be ensured by sampling according to the predictive mean $\mu_{t}$ of the surrogate, as $\exp\mu_{t}>0$ under mild assumptions. However, in practice, a non-zero value of $\beta$ helps to accelerate convergence in finite time by encouraging exploration. Corollary 3.1 is a direct application of Theorem C.3, proved in Appendix C.5, where we also contrast it with existing results [Oliveira et al., 2021].
Due to the reliance on GP-based approximations, a naive implementation of this approach would incur a cost of $\pazocal{O}(B^{3}t^{3})$ per stochastic gradient step, where $B$ is the mini-batch size, due to matrix inversions [Rasmussen and Williams, 2006]. However, for model spaces of moderate cardinality $|\pazocal{M}|=M$, we can keep compute costs linear with the number of optimization steps by applying recursive equations to evaluate the GP posterior mean and covariance over the model space (see Eq. 21 and 22), leading to a cost of $\pazocal{O}(B^{3}+MB^{2}+M^{2}B)=\pazocal{O}(M^{2}B)$ per step, as $B\ll M$, totaling $\pazocal{O}(TM^{2}B)$ over $T$ steps. Sparse approximations to GPs can further reduce this cost [Rasmussen and Williams, 2006, Gijsberts and Metta, 2013] to make it practical for larger model spaces. For our purposes, we implemented a diagonal Gaussian approximation, which makes the cost linear in the batch size and constant in $t$ via a mean-field approximation.
@@ 3.2 Categorical and neural probability mass functions
By Proposition˜D.1, we may represent probability distributions over the model space $\pazocal{M}$ by arbitrarily parametrized categorical distributions. A drawback of the surrogate is the need to maintain and update estimates over the entire model space, which can be impractical for spaces of very large cardinality, such as DAG discovery. Hence, we introduce two parametric alternatives.
Categorical:
Assume $|\pazocal{M}|=M\in{\mathbb{N}}$. Then, for $\psi\in{\mathbb{R}}^{M}$, the distribution over $\pazocal{M}$ is defined by $\smash{q_{\psi}(m):=(\sum_{j=1}^{M}\exp\psi_{j})^{-1}\sum_{i=1}^{M}\mathbb{I}[m_{i}=m]\exp\psi_{i}\,}$. The logit weights vector $\psi$ is unconstrained in ${\mathbb{R}}^{M}$ and can be jointly optimized with $\phi$ by gradient methods. Density evaluations and the entropy can be computed with memory cost $\pazocal{O}(|\pazocal{M}|)$.
Autoregressive:
If the model space is too large we may use a structured sample generation process which allows for the number of parameters to be smaller than cardinality of the model space i.e., $\dim(\psi)<M$. For instance, Germain et al. [2015] proposed an autoregressive parametrization for distributions over binary strings $\bm{s}\in\{0,1\}^{d_{s}}$ via the decomposition $\smash{p_{\psi}(\bm{s})=\prod_{i=1}^{d_{s}}p_{\psi}(s_{i}|s_{1},\dots,s_{i-1})}$. For each $\bm{s}$, we assign a unique $m\in\pazocal{M}$ and define $q_{\psi}(m):=p_{\psi}(\bm{s}(m))$. The conditional densities and sampling can be implemented via MADE, allowing us to map the entire model space with fewer parameters when $2^{d_{s}}\geq|\pazocal{M}|$. The same reasoning can be applied to a DAG via decomposition of its adjacency matrix. Details of MADE are in Appendix G.8, and for DAGs in Appendix G.7.
@@ 3.3 Estimation via Monte Carlo gradients
When $|\pazocal{M}|$ is too large to use a surrogate-based approach, or to even parameterize an entire vector of categorical weights in physical memory, we can employ neural-based methods that use gradient descent and estimation of the gradients of $\psi$ via Monte Carlo estimation of gradients (MCG) [Mohamed et al., 2020]. Using $\nabla_{\psi}q_{\psi}(m)=q_{\psi}(m)\nabla_{\psi}\log q_{\psi}(m)$, the gradient of the expectation in equation 9 with respect to $\psi$ is
@@
In practice, the variance of this estimator can be very high. However, techniques exist to reduce this variance [Paisley et al., 2012, Ranganath et al., 2014, Mohamed et al., 2020] for general applications. We use a control variate $\varsigma$ in the form
@@
where $g(\phi,\psi,\varsigma)=\mathbb{E}_{\bm{z}\sim\nu_{d_{\mathrm{max}}}}\left[\log h_{\phi}(\bm{z}|m)+\log q_{\psi}(m)-\log p(m)-\varsigma\right]$. We compute $\varsigma$ using the method described in Appendix E.1 (full description in Appendix E).
The benefit of using MCG for variational parameter estimation is the flexibility of choice for $q_{\psi}$. We compare two: (1) MCG of the logits of a standard categorical distribution, and (2) MCG of multi-layer perceptron weights that parameterize a configuration of the MADE neural autoregressive density estimator of Germain et al. [2015] (see Appendix G.7). When $|\pazocal{M}|$ is large, such implementations of $q_{\psi}$ permit an efficient approximate representation of the true model distribution.
3.4 Information-Limiting the optimization
The convergence of $\psi\rightarrow\psi^{*}$ is dependent on the convergence of $\phi\rightarrow\phi^{*}$, and optimal sample efficiency for the inference of $\phi$ is achieved when $\psi\approx\psi^{*}$. Intuitively, $q_{\phi}$ should focus primarily on the higher-probability models that contribute most to estimator variance, but discovering these models requires stable approximation of each $q_{\phi}(\bm{\theta}_{m}\mid m)$ to inform $\nabla_{\psi}$. This circular dependence motivates practical regulation of the optimization of $\psi$ when estimating $\nabla_{\psi}$ via Monte Carlo gradient estimates, addressing an instability similar to that discussed in the reinforcement learning literature [Schulman et al., 2017], but without modifying the objective. Our approach is to reduce the variance of the estimates of $q_{\psi}$ by bounding the information gain in the transition $q_{\psi_{t}}\to q_{\psi_{t+1}}$, which determines the step size, thereby stabilizing the optimization (detailed in Appendix E.2).
@@ 4 Related work
Conditional normalizing flows [Winkler et al., 2019, Durkan et al., 2019] have emerged as powerful tools for incorporating conditioning information. Existing methods use the context variable as a conditioning input, but fewer adapt the flow architecture itself. An exception is the transport‐based reversible jump MCMC method [Davies et al., 2023], which learns proposals for transdimensional moves, but does not readily allow its use as an inverse autoregressive flow [Kingma et al., 2016]. In contrast, we introduce an identity‐parameterized CoSMIC transformation without identity-map training. We bypass path-wise approximations to discrete distributions [Jang et al., 2017, Maddison et al., 2017], instead comparing Monte Carlo gradient estimation [Mohamed et al., 2020] with Bayesian optimization [Shahriari et al., 2016]. We adopt an information-based approach to scale gradient steps using “small steps,” inspired by reinforcement learning [Schulman et al., 2017]. Bayesian methods for model selection and optimization have advanced with black‐box variational inference [Ranganath et al., 2014, Titsias and Lázaro-Gredilla, 2014, Wingate and Weber, 2013] and flexible flows [Rezende and Mohamed, 2015, Papamakarios et al., 2017, Durkan et al., 2019]. Recent work in amortized Bayesian mixture models [Kucharsky and Burkner, 2025] shows amortization over multiple mixture components using conditional normalizing flows, but not for variable dimensions. Conversely, Li et al. [2020] introduces an architecture for learning imputation over transdimensional inputs, but lacks immediate application as a variational density. Our approach unifies transdimensional inference with flow‐based variational methods, bypassing the need for tailored dimension jumps and broadening applications.
@@ 5 Experiments
We present experiments involving synthetic and real data on two representative applications: robust variable selection and directed acyclic graphs. To evaluate the quality of the approximation $q_{\psi,\phi}(m,\bm{\theta}_{m})$ to the target distribution $\pi(m,\bm{\theta}_{m})$ for a relatively small $|\pazocal{M}|<2^{19}$ model space, we use the average negative log-likelihood (NLL) computed over a set of samples drawn from $\pi$ via a baseline sampling method, in this case reversible jump MCMC [Sisson, 2005]. Let $\{(m^{i},\bm{\theta}_{m}^{i})\}_{i=1}^{N}$ denote $N$ independent samples from $\pi(m,\bm{\theta}_{m})$. The average NLL corresponds to the cross-entropy $H(\pi,q_{\psi,\phi})$ between $\pi$ and $q_{\psi,\phi}$, which quantifies the expected number of bits needed to encode samples from $\pi$ using $q_{\psi,\phi}$, and is defined as $\smash{\text{NLL}=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi,\phi}(m^{i},\bm{\theta}_{m}^{i})}$. Comparison of VTI DAG inference quality with baseline frequentist and Bayesian approaches use standard metrics [Kummerfeld and Rix, 2019].